OpenStudy (anonymous):

4/3 times 3.14 times 3.1 cubed?

OpenStudy (anonymous):

\[\frac{ 4 }{ 3 }\times \pi \times 3.1^{3}\] Is that it?

OpenStudy (anonymous):

Yes!

OpenStudy (anonymous):

\[3.1^{3}=29.791 \times \pi=93.591\times \frac{ 4 }{ 3 } = 124.788\]